OpenStudy (anonymous):

evaluate the integral x/(sqrt(x^2+x+2) dx

OpenStudy (anonymous):

so its like this?: \[\int\limits_{ }^{ }\frac{ x }{ \sqrt{x^2 +x+2} }dx\]

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Hmm... Tough one...

OpenStudy (anonymous):

Alright I solved it but the guy isn't on so I guess i'll wait.